OpenStudy (anonymous):

PLEASE HELP! WILL FAN AND MEDAL! :):) Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 − 4x + y2 + 8y = −4

OpenStudy (amitrashok):

lol u do it it is easy

OpenStudy (anonymous):

@jim_thompson5910 sorry.. i didn't realize i closed it

OpenStudy (anonymous):

this is all i got... The general equation of a circle is: x2+y2+2gx+2fy+c=0....where the center is: (−g,−f) so if you subtract 4 from both sides of the equation you get: x2+y2−4x+8y+4=0 2g is -4 and 2f is 8 so the center is (-g,-f) which is (2,-4)

OpenStudy (anonymous):

i remember most of it.. but i know im missing something in the equation..

jimthompson5910 (jim_thompson5910):

if we just focus on x^2-4x, what do we need to complete the square?

jimthompson5910 (jim_thompson5910):

example: say we had x^2 + 12x to complete the square, we take half of 12 (the x coefficient) to get 6 then we square 6 to get 36. This is added and subtracted onto x^2+12x to get `x^2+12x+36-36` notice how the `x^2+12x+36` factors to `(x+6)^2` so `x^2+12x` turns into `x^2+12x+36-36` which becomes `(x+6)^2-36`

OpenStudy (anonymous):

oh.. i think ive been doing this all wrong...

OpenStudy (anonymous):

i did... (x² − 4x) + (y² + 8y) = −4 (x² - 4x + 2²) + (y² + 8y + 4²) = -4 + 2² + 4² (x-2)² + (y+4)² = 4² center (2,-4)

jimthompson5910 (jim_thompson5910):

What you posted above is correct, but this is how I'm familiar with this type of problem.

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

you have the correct steps, that leads to the radius being r = 4

jimthompson5910 (jim_thompson5910):

you also have the correct center

OpenStudy (anonymous):

oh.. well.. how did you get 4?

jimthompson5910 (jim_thompson5910):

see attached

jimthompson5910 (jim_thompson5910):

Your last equation is in the form \[\Large (x-h)^2 + (y-k)^2 = r^2\] the right side, which is 4^2 is equal to r^2 r^2 = 4^2 r = 4

OpenStudy (anonymous):

oh wow.. now i feel really stupid haha

jimthompson5910 (jim_thompson5910):

nah, you just overlooked it

OpenStudy (anonymous):

thank you for your help! i really appreciate it :)

jimthompson5910 (jim_thompson5910):

you're welcome